# Simple homotopy types of even dimensional manifolds

### v2

@@ 10.5. Divisibility and lower bounds for kernel groups
The aim of this section will be to establish divisibility results for $|D(\mathbb{Z}C_{m})|$ and $\operatorname{odd}(|\widetilde{V}_{m}|)$. These results are necessary for determining the involution on $D(\mathbb{Z}C_{m})$ in an analogous way to how divisibility results for class numbers $h_{m}$ were necessary for determining the involution on $C(\mathbb{Z}[\zeta_{m}])$ (see Section 10.3). The results on $\operatorname{odd}(|\widetilde{V}_{m}|)$ are motivated by Proposition 10.17.
@@
We will now find conditions on $m\geq 2$ square-free for which $\operatorname{odd}(|\widetilde{V}_{m}|)\neq 1$. Our strategy is motivated by the bounds $d$ such that $d\mid|D(\mathbb{Z}C_{m})|$ which were obtained by Cassou-Noguès in [CN72, CN74]. In particular, our argument shows that these bounds actually give factors of $|\widetilde{V}_{m}|$. Recall that $\widetilde{V}_{m}=\operatorname{coker}(\Psi_{m}^{+})$, where
@@
The proofs of (ii) and (iii) are analogous. We will prove (iii) only as it is more complicated. First note that $\mathbb{Q}(\lambda_{m})/\mathbb{Q}$ is a Galois extension and $p$ is unramified in $\mathbb{Q}(\lambda_{m})/\mathbb{Q}$ since it is unramified in $\mathbb{Q}(\zeta_{m})/\mathbb{Q}$ (see, for example, [Wa97, Lemma 15.48]). This implies that $p\cdot\mathbb{Z}[\lambda_{m}]=\mathcal{P}_{1}\cdots\mathcal{P}_{g}$ for some $g\geq 1$, where the $\mathcal{P}_{i}\subseteq\mathbb{Z}[\lambda_{m}]$ are distinct prime ideals. The $\mathcal{P}_{i}$ coincide under the Galois action and the $\mathbb{Z}[\lambda_{m}]/\mathcal{P}_{i}$ all coincide with the splitting field $\mathbb{F}_{p}(\lambda_{m})$ and so
@@
Let $f=[\mathbb{F}_{p}(\lambda_{m}):\mathbb{F}_{p}]$. Since $\operatorname{Gal}(\mathbb{F}_{p}(\lambda_{m})/\mathbb{F}_{p})$ is generated by the Frobenius element $\operatorname{Frob}_{p}\colon x\mapsto x^{p}$, we get that $f$ is the smallest positive integer such that $\operatorname{Frob}_{p}^{f}=\operatorname{Id}_{\mathbb{F}_{p}(\lambda_{m})}$. Note that $\operatorname{Frob}_{p}^{f}(\lambda_{m})=\zeta_{m}^{p^{f}}+\zeta_{m}^{-p^{f}}$. This implies that $\operatorname{Frob}_{p}^{f}=\operatorname{Id}_{\mathbb{F}_{p}(\lambda_{m})}$ if and only if $\zeta_{m}^{p^{f}}=\zeta_{m}^{\pm 1}$ and so $f$ is the order of $p$ in $(\mathbb{Z}/m)^{\times}/\{\pm 1\}$. Since $[\mathbb{Q}(\lambda_{m}):\mathbb{Q}]=\varphi(m)/2$, we have that $|\mathbb{F}_{p}[\lambda_{m}]|=p^{\varphi(m)/2}$. Since $|\mathbb{F}_{p}(\lambda_{m})^{g}|=p^{fg}$, this gives that $g=\varphi(m)/2f$. Hence we have $|\mathbb{F}_{p}[\lambda_{m}]^{\times}|=|\mathbb{F}_{p}(\lambda_{m})^{\times}|^{g}=(p^{f}-1)^{g}$. Note that $f=f_{p}$ if and only if $f_{p}$ is odd, and otherwise $f=f_{p}/2$.
@@
We will now use Lemma 10.21 to obtain bounds on $|\widetilde{V}_{pq}|$ for $p,q$ distinct odd primes, and for $|\widetilde{V}_{2p}|$, where $p$ is an odd prime respectively.
@@
$c_{pq}=\begin{cases}\frac{1}{2pq}(p^{\frac{f_{p}}{2}}+1)^{g_{p}}(q^{\frac{f_{q}}{2}}+1)^{g_{q}},&\text{if $f_{p},f_{q}$ are even}\\ \frac{1}{2pq}(p^{\frac{f_{p}}{2}}+1)^{g_{p}}(q^{f_{q}}-1)^{\frac{g_{q}}{2}},&\text{if $f_{p}$ is even and $f_{q}$ is odd}\\ \frac{1}{2pq}(p^{f_{p}}-1)^{\frac{g_{p}}{2}}(q^{\frac{f_{q}}{2}}+1)^{g_{q}},&\text{if $f_{p}$ is odd and $f_{q}$ is even}\\ \frac{1}{2pq}(p^{f_{p}}-1)^{\frac{g_{p}}{2}}(q^{f_{q}}-1)^{\frac{g_{q}}{2}},&\text{if $f_{p},f_{q}$ are odd.}\end{cases}$
@@
where $1\in\mathbb{Z}/pq$ maps to $\Psi_{pq}^{+}(\zeta_{pq})$ and $21\in\mathbb{Z}/$ 2 maps to $\Psi_{pq}^{+}(1-\zeta_{pq})$. In particular, this implies that $|\widetilde{V}_{pq}|$ is divisible by $\frac{1}{2pq}\cdot|\frac{\mathbb{F}_{p}[\zeta_{q}]^{\times}}{\mathbb{F}_{p}[\lambda_{q}]^{\times}}|\cdot|\frac{\mathbb{F}_{q}[\zeta_{p}]^{\times}}{\mathbb{F}_{q}[\lambda_{p}]^{\times}}|$. The result now follows from Lemma 10.21. ∎
@@
[…]
@@ 11. Proof of main results on the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$
The aim of this section will be to prove the following three theorems. As we saw previously, these theorems are required to prove Theorems 5.14, 5.15, LABEL:, and 5.17 respectively.
@@
    |$|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,17,19\}$; and
@@
    |$|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|\to\infty$ super-exponentially in $m$.
Theorem 11.2.
Let $m\geq 2$ be a square-free integer. Then
  1. (i)
    |$|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$; and
  2. (ii)
    |$|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\to\infty$ super-exponentially in $m$.
@@
    |$|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}/\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|=1$ for infinitely many $m$; and
@@
    $\sup_{n\leq m}|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}/\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\to\infty$ exponentially in $m$.
11.1. Proof of Theorems 11.1 and 11.2
The proofs of Theorems 11.1 and 11.2 are best handled together since
$\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}\leq\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}.$
In particular, it suffices to prove the following four statements.
@@
    |$|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\to\infty$ super-exponentially in $m$.
@@
    If $m\not\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$, then |$|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\neq 1$.
@@
    If $m\in\{2,3,5,6,7,10,11,13,14,17,19\}$, then |$|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|=1$.
@@
    If $m\in\{15,29\}$, then |$|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|=1$ and |$|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|\neq 1$.
To see this, note that (A1) coincides with Theorem 11.2 (ii) and implies Theorem 11.1 (ii). The forwards direction of Theorem 11.1 (i) is implied by (A2) and (A4), and the backwards direction coincides with (A3). The forwards direction of Theorem 11.2 (i) coincides with (A2) and the backwards direction is implied by (A3) and (A4).
@@
where $D(\mathbb{Z}C_{m})$ has the induced involution, and each $C(\mathbb{Z}[\zeta_{d}])$ has the involution induced by conjugation. Each of statements (A1)–(A4) will be proven via the following lemma.
@@
    |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|\leq|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|$;
@@
    if |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|\neq 1$ or |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|\neq 1$, then |$|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\neq 1$ ((((and so |$|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|\neq 1)$;
@@
    if |$|\{x\in D(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|=1$ and $\prod_{d\mid m}|\{x\in C(\mathbb{Z}[\zeta_{d}])\mid\overline{x}=-x\}|=1$, then |$|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|=1$ ((((and so |$|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|=1)$.
@@
$\qquad\qquad\to\bigoplus_{d\mid m}\{x\in C(\mathbb{Z}[\zeta_{d}])\mid x=-\overline{x}\}\vspace{-5mm}$
@@
We will now proceed to prove each of statements (A1)-(A4). We will begin with the following which, by Lemma 11.4 (ii), implies (A1).
@@
We have that |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|\to\infty$ super-exponentially in $m$.
@@
It was shown in Proposition 10.4 that $\operatorname{odd}(h_{m}^{-})$ divides |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|$. The result now follows from that fact that, by Proposition 10.8, $\operatorname{odd}(h_{m}^{-})\to\infty$ super-exponentially in $m$. ∎
We will now prove (A2). Our approach will be to use Lemma 11.4 (iii). In particular, we will begin by classifying the $m\geq 2$ square-free for which |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$. We will then determine the subset of these values for which |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$.
@@
The complete list of $m\geq 2$ square-free for which |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$ is as follows:
@@
First note that $h_{m}^{-}=1$ implies |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$. On the other hand, it follows from Proposition 10.4 that $\operatorname{odd}(h_{m}^{-})\neq 1$ implies |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|\neq 1$. In Proposition 10.7 (i), it is shown that the $m\geq 2$ square-free for which $\operatorname{odd}(h_{m}^{-})=1$ and $h_{m}^{-}\neq 1$ are precisely the $m\in S$, where $S=\{29,39,58,65,78,130\}$. It remains to determine for which $m\in S$ we have |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$.
@@
If $m=29$ or $68$, then $C(\mathbb{Z}[\zeta_{m}])\cong(\mathbb{Z}/2)^{3}$ and $02\cdot C(\mathbb{Z}[\zeta_{m}])=$ 0 [Wa97, p. 412]. If $m=39$ or $78$, then $h_{m}=2$ and $C(\mathbb{Z}[\zeta_{m}])\cong\mathbb{Z}/2$ and $02\cdot C(\mathbb{Z}[\zeta_{m}])=$ 0 [Wa97, p. 412]. If $m=65$ or $130$, then [Ho93, Proposition 1 (iv)] gives that $C(\mathbb{Z}[\zeta_{m}])\cong(\mathbb{Z}/2)^{2}\times(\mathbb{Z}/4)^{2}$ and $2\cdot C(\mathbb{Z}[\zeta_{m}])\cong(\mathbb{Z}/2)^{2}$.
Hence we have shown that, for $m\geq 2$ square-free, |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$ if and only if $h_{m}^{-}=1$ or $m\in\{29,39,58,78\}$. The result now follows by Proposition 10.5 (i). ∎
@@
The complete list of $m\geq 2$ square-free for which |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$ and |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$ is as follows:
@@
Before turning to the proof, we will begin by recalling that Cassou-Noguès determined the integers $m\geq 2$ for which $D(\mathbb{Z}C_{m})=0$ [CN74, Theorème 1] (see also [CN72]). The following can be deduced by comparing this with Proposition 11.6.
@@
Let $m\geq 2$ be a square-free integer such that |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$. Then $D(\mathbb{Z}C_{m})=0$ if and only if $m$ is prime or $m=2p$ where $p\in\{3,5,7\}$.

### v3

@@ 10.5. Divisibility and lower bounds for kernel groups
The aim of this section is to establish divisibility results for $|D(\mathbb{Z}C_{m})|$ and $\operatorname{odd}(|\widetilde{V}_{m}|)$. These results are necessary for determining the involution on $D(\mathbb{Z}C_{m})$ in an analogous way to how divisibility results for class numbers $h_{m}$ were necessary for determining the involution on $C(\mathbb{Z}[\zeta_{m}])$ (see Section 10.3). The results on $\operatorname{odd}(|\widetilde{V}_{m}|)$ are motivated by Proposition 10.18.
@@
We now find conditions on $m\geq 2$ square-free for which $\operatorname{odd}(|\widetilde{V}_{m}|)\neq 1$. Our strategy is motivated by the bounds $d$ such that $d\mid|D(\mathbb{Z}C_{m})|$ which were obtained by Cassou-Noguès in [CN72, CN74]. In particular, our argument shows that these bounds actually give factors of $|\widetilde{V}_{m}|$. Recall that $\widetilde{V}_{m}=\operatorname{coker}(\Psi_{m}^{+})$, where
@@
The proofs of (ii) and (iii) are analogous. We prove (iii) only as it is more complicated. Firstly, $\mathbb{Q}(\lambda_{m})/\mathbb{Q}$ is a Galois extension and $p$ is unramified in $\mathbb{Q}(\lambda_{m})/\mathbb{Q}$ since it is unramified in $\mathbb{Q}(\zeta_{m})/\mathbb{Q}$ (see, for example, [Wa97, Lemma 15.48]). This implies that $p\cdot\mathbb{Z}[\lambda_{m}]=\mathcal{P}_{1}\cdots\mathcal{P}_{g}$ for some $g\geq 1$, where the $\mathcal{P}_{i}\subseteq\mathbb{Z}[\lambda_{m}]$ are distinct prime ideals. The $\mathcal{P}_{i}$ coincide under the Galois action and the $\mathbb{Z}[\lambda_{m}]/\mathcal{P}_{i}$ all coincide with the splitting field $\mathbb{F}_{p}(\lambda_{m})$ and so
@@
Let $f=[\mathbb{F}_{p}(\lambda_{m}):\mathbb{F}_{p}]$. Since $\operatorname{Gal}(\mathbb{F}_{p}(\lambda_{m})/\mathbb{F}_{p})$ is generated by the Frobenius element $\operatorname{Frob}_{p}\colon x\mapsto x^{p}$, we get that $f$ is the smallest positive integer such that $\operatorname{Frob}_{p}^{f}=\operatorname{Id}_{\mathbb{F}_{p}(\lambda_{m})}$. Note that $\operatorname{Frob}_{p}^{f}(\lambda_{m})=\zeta_{m}^{p^{f}}+\zeta_{m}^{-p^{f}}$. This implies that $\operatorname{Frob}_{p}^{f}=\operatorname{Id}_{\mathbb{F}_{p}(\lambda_{m})}$ if and only if $\zeta_{m}^{p^{f}}=\zeta_{m}^{\pm 1}$ and so $f$ is the order of $p$ in $(\mathbb{Z}/m)^{\times}/\{\pm 1\}$. Since $\mathbb{Q}(\lambda_{m}):\mathbb{Q}]=\varphi(m)/2$, we have that $|\mathbb{F}_{p}[\lambda_{m}]|=p^{\varphi(m)/2}$. Since $|\mathbb{F}_{p}(\lambda_{m})^{g}|=p^{fg}$, this gives that $g=\varphi(m)/2f$. Hence we have $|\mathbb{F}_{p}[\lambda_{m}]^{\times}|=|\mathbb{F}_{p}(\lambda_{m})^{\times}|^{g}=(p^{f}-1)^{g}$. Note that $f=f_{p}$ if and only if $f_{p}$ is odd, and otherwise $f=f_{p}/2$.
@@
We now use Lemma 10.22 to obtain bounds on $|\widetilde{V}_{pq}|$ for $p,q$ distinct odd primes, and for $|\widetilde{V}_{2p}|$, where $p$ is an odd prime respectively.
@@
[…]
@@
where $1\in\mathbb{Z}/pq$ maps to $\Psi_{pq}^{+}(\zeta_{pq})$ and $1\in\mathbb{Z}/2$ maps to $\Psi_{pq}^{+}(1-\zeta_{pq})$. In particular, this implies that $|\widetilde{V}_{pq}|$ is divisible by $\frac{1}{2pq}\cdot|\frac{\mathbb{F}_{p}[\zeta_{q}]^{\times}}{\mathbb{F}_{p}[\lambda_{q}]^{\times}}|\cdot|\frac{\mathbb{F}_{q}[\zeta_{p}]^{\times}}{\mathbb{F}_{q}[\lambda_{p}]^{\times}}|$. The result now follows from Lemma 10.22. ∎
@@
$\mbox{\footnotesize$\displaystyle\frac{1}{p}$}\cdot\Big|\mbox{\footnotesize$\displaystyle\frac{\mathbb{F}_{p}[\zeta_{2}]^{\times}}{\mathbb{F}_{p}[\lambda_{2}]^{\times}}$}\Big|\cdot\Big|\mbox{\footnotesize$\displaystyle\frac{\mathbb{F}_{2}[\zeta_{p}]^{\times}}{\mathbb{F}_{2}[\lambda_{p}]^{\times}}$}\Big|=\mbox{\footnotesize$\displaystyle\frac{1}{p}$}\cdot\Big|\mbox{\footnotesize$\displaystyle\frac{\mathbb{F}_{2}[\zeta_{p}]^{\times}}{\mathbb{F}_{2}[\lambda_{p}]^{\times}}$}\Big|.$
@@ 11. Proof of main results on the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$
The aim of this section is to prove the following theorems. Theorem 11.1 coincides with Theorem 1.7 and Theorem 11.2 is a more detailed version of Theorem 1.8. As we saw previously, these theorems are required to prove Theorems 5.14 and 5.15 respectively.
@@
    $|\{y\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{y}=-y\}|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,17,19\}$;
@@
    $|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$; and
  3. (iii)
    $|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\to\infty$ super-exponentially in $m$, and hence we also have that $|\{y\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{y}=-y\}|\to\infty$ super-exponentially in $m$.
@@
    $|\{y\in\widetilde{K}_{0}(\mathbb{Z}C_{p^{n}})\mid\overline{y}=-y\}/\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{p^{n}})\}|=1$ if $n\geq 1$ and $p\leq 509$ is an odd prime such that $h_{p}$ is odd (e.g. $p=3$);
@@
    The sequence $|\{y\in\widetilde{K}_{0}(\mathbb{Z}C_{2^{n}})\mid\overline{y}=-y\}/\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{2^{n}})\}|$ is unbounded as $n\to\infty$. Thus, $\sup_{k\leq m}|\{y\in\widetilde{K}_{0}(\mathbb{Z}C_{k})\mid\overline{y}=-y\}/\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{k})\}|\to\infty$ exponentially in $m$.
11.1. Proof of Theorem 11.1
In light of the fact that
$\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}\leq\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\},$
it suffices to prove the following four statements.
@@
    $|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\to\infty$ super-exponentially in $m$.
@@
    If $m\not\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$, then $|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\neq 1$.
@@
    If $m\in\{2,3,5,6,7,10,11,13,14,17,19\}$, then $|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|=1$.
@@
    If $m\in\{15,29\}$, then $|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|=1$ and $|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|\neq 1$.
To see this, note that (A1) implies Theorem 11.1 (iii). The forwards direction of Theorem 11.1 (iii) is implied by (A2) and (A4), and the backwards direction coincides with (A3). The forwards direction of Theorem 11.1 (ii) coincides with (A2) and the backwards direction is implied by (A3) and (A4).
@@
where $D(\mathbb{Z}C_{m})$ has the induced involution, and each $C(\mathbb{Z}[\zeta_{d}])$ has the involution induced by conjugation. Each of statements (A1)–(A4) are proven via the following lemma.
@@
    $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|\leq|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|$;
@@
    if $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|\neq 1$ or $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|\neq 1$, then $|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|\neq 1$ ((and so $|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|\neq 1)$;
@@
    if $|\{x\in D(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|=1$ and $\prod_{d\mid m}|\{x\in C(\mathbb{Z}[\zeta_{d}])\mid\overline{x}=-x\}|=1$, then $|\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{x}=-x\}|=1$ ((and so $|\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}|=1)$.
@@
$\qquad\qquad\to\bigoplus_{d\mid m}\{x\in C(\mathbb{Z}[\zeta_{d}])\mid x=-\overline{x}\}\vskip-14.22636pt$
@@
We now proceed to prove each of statements (A1)-(A4). We begin with the following which, by Lemma 11.3 (ii), implies (A1).
@@
We have that $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|\to\infty$ super-exponentially in $m$.
@@
It was shown in Proposition 10.5 that $\operatorname{odd}(h_{m}^{-})$ divides $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|$. The result now follows from that fact that, by Proposition 10.9, $\operatorname{odd}(h_{m}^{-})\to\infty$ super-exponentially in $m$. ∎
We now prove (A2). Our approach is to use Lemma 11.3 (iii). In particular, we begin by classifying the $m\geq 2$ square-free for which $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$. We then determine the subset of these values for which $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$.
@@
The complete list of $m\geq 2$ square-free for which $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$ is as follows:
@@
Firstly, $h_{m}^{-}=1$ implies $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$. On the other hand, it follows from Proposition 10.5 that $\operatorname{odd}(h_{m}^{-})\neq 1$ implies $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|\neq 1$. In Proposition 10.8 (i), it is shown that the $m\geq 2$ square-free for which $\operatorname{odd}(h_{m}^{-})=1$ and $h_{m}^{-}\neq 1$ are precisely the $m\in S$, where $S=\{29,39,58,65,78,130\}$. It remains to determine for which $m\in S$ we have $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$.
@@
If $m=29$ or $68$, then $C(\mathbb{Z}[\zeta_{m}])\cong(\mathbb{Z}/2)^{3}$ and $2\cdot C(\mathbb{Z}[\zeta_{m}])=0$ [Wa97, p. 412]. If $m=39$ or $78$, then $h_{m}=2$ and $C(\mathbb{Z}[\zeta_{m}])\cong\mathbb{Z}/2$ and $2\cdot C(\mathbb{Z}[\zeta_{m}])=0$ [Wa97, p. 412]. If $m=65$ or $130$, then [Ho93, Proposition 1 (iv)] gives that $C(\mathbb{Z}[\zeta_{m}])\cong(\mathbb{Z}/2)^{2}\times(\mathbb{Z}/4)^{2}$ and $2\cdot C(\mathbb{Z}[\zeta_{m}])\cong(\mathbb{Z}/2)^{2}$.
Hence we have shown that, for $m\geq 2$ square-free, $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$ if and only if $h_{m}^{-}=1$ or $m\in\{29,39,58,78\}$. The result now follows by Proposition 10.6 (i). ∎
@@
The complete list of $m\geq 2$ square-free for which $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$ and $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$ is as follows:
@@
Before turning to the proof, we begin by recalling that Cassou-Noguès determined the integers $m\geq 2$ for which $D(\mathbb{Z}C_{m})=0$ [CN74, Theorème 1] (see also [CN72]). The following can be deduced by comparing this with Proposition 11.5.
@@
Let $m\geq 2$ be a square-free integer such that $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$. Then $D(\mathbb{Z}C_{m})=0$ if and only if $m$ is prime or $m=2p$ where $p\in\{3,5,7\}$.